OpenStudy (anonymous):

What is the slope of the graph of 7x + 14y = -9? A. -2 B. -½ C. ½ D. 2

OpenStudy (anonymous):

.....

OpenStudy (tfht):

ya solve for y so to do that you need to get y alone by subtracting 7x from both sides to get 14y=-7y-9

OpenStudy (tfht):

then you divide both sides by 14 to get y=-1/2x-9/14. so do you know what y=mx+b means?

OpenStudy (tfht):

m=the slope and b=the y coordinate. so if m=slope and -1/2 is in m's spot, then what's the slope?

OpenStudy (anonymous):

Elaborate? Yes I can. I was just confused

OpenStudy (anonymous):

Okay.. Thank you

OpenStudy (anonymous):

@tfht thank you so much

OpenStudy (anonymous):

for explaining

OpenStudy (anonymous):

So would it be -2?

OpenStudy (tfht):

no what is in front of x?